OpenStudy (heart_offire):

Find an equation of the line that passes through the points (1,2) and (2,3). A) y = x - 1 B) y = x + 1 C) y = x - 2 D) y = 2x - 1 @pooja195 i dont know how to start on this one

pooja195 (pooja195):

Start by finding the slope Use the slope formula \[ \huge~\rm~\bf\color{blue}{ \frac{y_2-y_1}{x_2-x_1}=slope} \]

OpenStudy (heart_offire):

1/-1

pooja195 (pooja195):

\[\huge~\rm~\bf\color{blue}{ \frac{3-2}{2-1}=slope}\]

OpenStudy (heart_offire):

do i insert the slope into y=mx+b

pooja195 (pooja195):

you have the incorrect slope. 3-2=? 2-1=?

OpenStudy (heart_offire):

3-2=1 and 2-1= -1

pooja195 (pooja195):

2-1 does not equal -1

OpenStudy (heart_offire):

1 sorry

pooja195 (pooja195):

Good 1/1=1 soo... our slope is just x y=mx+b m=1 now we have: y=1x+b To find the b we need to plug in one of the points lets do (1,2) y=1x+b Plug in your coordinates x=1 y=2 Plug in: 2=1(1)+b solve for b

OpenStudy (heart_offire):

umm b=1?

pooja195 (pooja195):

Perfect! :) y=1x+b now plug in the 1 in place of b what's our answer? (final answeR)

OpenStudy (heart_offire):

so that answe would be B

pooja195 (pooja195):

Good job :-)

OpenStudy (heart_offire):

thanks:)

pooja195 (pooja195):

yw ^_^